# Faux Polyglot: A Study on Information Disparity in Multilingual Large Language Models

### v1

@@ Abstract
With Retrieval Augmented Generation (RAG), Large Language Models (LLMs) are playing a pivotal role in information search and are being adopted globally. Although the multilingual capability of LLMs offers new opportunities to bridge the language barrier, do these capabilities translate into real-life scenarios where linguistic divide and knowledge conflicts between multilingual sources are known occurrences? In this paper, we studied LLM’s linguistic preference in a RAG-based information search setting. We found that LLMs displayed systemic bias towards information in the same language as the query language in both information retrieval and answer generation. Furthermore, in scenarios where there is little information in the language of the query, LLMs prefer documents in high-resource languages, reinforcing the dominant views. Such bias exists for both factual and opinion-based queries. Our results highlight the linguistic divide within multilingual LLMs in information search systems. The seemingly beneficial multilingual capability of LLMs may backfire on information parity by reinforcing language-specific information cocoons or filter bubbles further marginalizing low-resource views.
@@
Figure 1: The figure is based on real-world interactions with GPT-4. The figure shows users interacting with GPT-4 OpenAI (2024) using 4 languages (en, ar, hi, zh). When the contexts are available in 3 (en,zh,hi) of the 4 languages, the queries in different languages resulted in drastically different answers for each language. We see 2 effects in this figure: 1) Information cocoon – information presented to the users of a particular language is from their own language and 2) Cultural Dominance – information asked from a lower resource language where the information is not available in, results in LLM choosing dominant or highest resource language document as the information source leading to reinforcement of dominant views.
1 Introduction
@@
Large language models (LLMs) have played a pivotal role in search technologies such as ChatGPT [^1] https://chatgpt.com/ and Bing Copilot [^2] https://www.bing.com/chat. With Retrieval Augmented Generation (RAG), LLMs can leverage external knowledge that is outside of their training data and respond to questions about ongoing events Lewis et al. (2021). Such search technologies have significantly impacted how people seek and access information worldwide. More importantly, the premise of LLM’s multilingual capability Qin et al. (2024a), supporting over 100 languages and multilingual understanding, offers new opportunities to break the language-specific information cocoon and provide more equitable information access for people with diverse linguistic preferences, cultural backgrounds, and geographical locations.
@@
However, achieving information parity is challenging because of the intrinsic complexity in how information is created and represented in different linguistic contexts. This complexity stems from a broad spectrum of cultural nuances, regional influences, and historical narratives that shape how information is consumed and communicated White (1990); Edelman (1985). In addition, each language and cultural context also brings its own set of values and perspectives when interpreting an event of interest, which can significantly influence the representation of facts and narratives Foucault (1980).
@@
This challenge is exacerbated by the linguistic divide, the gap in terms of representation between different languages on the internet Li et al. (2024a); Mehdad et al. (2012). Not all facts and viewpoints are presented equally. Dominant languages, such as English, have a disproportionate influence on content creation and dissemination Holborow (1996). Such divides can lead to linguistic imperialism, where information in higher-resource languages is more accessible and frequently amplified, potentially overshadowing or distorting narratives in low-resource languages, particularly when differing views are presented Phillipson (2018).
@@
As a result, when responding to user’s queries, LLMs need to retrieve meaningful information from diverse multilingual sources and synthesize diverse perspectives and often conflicting knowledge residing in multilingual contexts. For example, when an information seeker asks for an event of conflict or topic with multiple perspectives in different regions, for instance, an ongoing war between two countries with distinct languages, LLMs should be able to present viewpoints and facts from both sides regardless of the query language.
@@
Recent evidence indicates information disparity in LLMs’ behavior. Individuals asking the same query in different languages may get different answers. This difference manifests in the quality of the responses Boughorbel and Hawasly (2023); Jin et al. (2024a), consistency Dong et al. (2024), and dispute resolution Li et al. (2023). As illustrated in Fig.1, current models like GPT-4 struggle with information disparity in multilingual settings. The narratives generated by LLMs regarding a conflict can vary significantly depending on the language of the query and the language of reference documents.
Therefore, it is crucial to understand how the linguistic divide affects information disparity in the context of multilingual LLMs. In RAG, these biases can potentially skew both the retrieval and generation processes, favoring information from certain languages over others. Such biases could undermine the very goal of using multilingual LLMs to democratize access to information and support diverse information-seeking globally. If unchecked, these biases could reinforce existing cultural dominance and create an information cocoon, potentially alienating non-dominant language speakers and disproportionately perpetuating dominant views.
This study aims to understand information parity in multilingual LLMs, especially when conflicting perspectives and facts are presented in different languages. Specifically, we ask the following research questions,
@@
To answer our research questions, we proposed a new multilingual synthetic dataset with 170 documents in 5 languages providing information about two fictional events. The dataset represents a real-world information environment where diverse, sometimes conflicting viewpoints and factual information are represented in different languages. Using the dataset as source references, we studied 13 multilingual LLMs with 7 query types and found,
  • Current multilingual LLMs have a systematic preference towards documents in the Native Language of the query.
  • Such preference is consistent across both factual and opinion-based queries in the retrieval phase.
  • Factual queries specifically targeting information in Foreign Languages can successfully retrieve information from Foreign Documents in the generation phase.
  • When there is no relevant document in the language of the query, the current LLMs prefer documents in high-resource languages over low-resource languages.
Our results indicate that while current multilingual LLMs claim to create more equitable information access, the linguistic divide in those LLMs may create a language-specific information cocoon and amplify perspectives and facts in the dominant language, especially when handling knowledge conflicts in RAG-based search systems.
@@ 2 Experiments
There are two types of LLM-powered search systems: Retrieval Augmented Generation (RAG) and Direct Generation. In Direct Generation, the LLM relies on its parametric memory to respond to the user’s query. In this paper, we focus on RAG Lewis et al. (2021), where LLMs primarily rely on external knowledge presented in the context to answer user information queries. We made this choice for the following reasons: 1) most LLMs are closed source with no access to their parametric memory, and 2) most popular information systems, such as Bing Copilot, Perplexity.ai  [^3] https://www.perplexity.ai/, etc., use RAG to keep their search system updated with the most recent information.
@@
Retrieval Augmented Generation consists of two phases: retrieval and generation. In the retrieval phase, the system retrieves a ranked list of documents based on cosine similarity scores between query embedding and document embedding. In the generation phase, the top documents are used as LLM’s context to generate the response. In this study, we separated the two phases to independently study the effect of language when multilingual LLMs answer user queries with RAG.
Isolating LLM’s Parametric Memory
To study LLM’s information preference in RAG, we need to separate the effects of parametric memory from in-context non-parametric memory. We created a synthetic dataset with fictional entities that were not in the LLM’s pre-training data.
@@ 2.1 Multilingual Fictional Dataset
Figure 2: The figure illustrates our fictional dataset creation process. We manually generated core facts- short, individual facts forming a story. Using GPT-4-turboOpenAI (2024), we elaborated on these core facts, ensuring no contradictions. Different subsets of these extended core facts were used to prompt the LLM to create News and Story documents. This process is repeated to produce both the original and alternate document sets. Finally, we use Google Translate via translate-shell to translate the documents into German, Chinese, Arabic, and Hindi. Details of prompts are in Appendix.
The overarching goal of our dataset is to mimic a real-world multilingual information environment where different, sometimes conflicting, facts and perspectives are represented in different languages. The dataset consists of 170 documents that were later used in our experiment to answer user queries.
Our dataset supports two common information queries: facts and opinions. We considered the following scenarios for each type of query:
Discrepancies in facts:
  • Shared Facts: same facts shared across different languages.
Diverse opinions and perspectives:
@@
We selected English (en), Hindi (hi), German (de), Arabic (ar), and Chinese (zh) as our target languages. These languages represent 4 different scripts, different cultures, and a mix of high-resource (en, de, zh) and low-resource languages (ar, hi). The chosen languages account for around 4 billion active speakers around the world.
@@ 2.1.1 Premise setup
The festival premise is designed for factual queries, which contain information about a fictional festival targeting factual queries. There are versions: original and alternate – each with some shared, contradictory, and exclusive facts. Here is an example with contradictory facts,
@@
The war premise is designed for opinion-based queries. It contains information about a fictional war taking place between two fictional regions speaking different languages. This premise is designed to represent an information environment with diverse opinions and perspectives. Similarly, there are two document sets in the war: original and alternate – each taking the perspective of one side in the war. Example perspectives are
@@
2.1.2 Synthetic Dataset Creation
Core facts:
Extended story facts:
Final Documents:
Manipulation Check:
2.1.3 Multilingual Dataset Creation
@@ 2.2 Experiment Design
We queried popular multilingual retrieval models and multilingual LLMs with different types of user queries to analyze the preference of models when presented with conflicting multilingual information in retrieval and generation.
@@ 2.2.1 Diverse Information Queries
We curated a diverse set of queries to mimic real-world multilingual information-seeking behaviors. We covered two types of information queries: factual and opinion-based. Each query has versions in 5 languages: English(en), Hindi(hi), Chinese(zh), Arabic(ar), and German(de) aligned with the languages used in the dataset.
@@
Based on the core facts, we created 27 factual queries consisting of 6 broad, 9 specific-contradiction, 9 specific-exclusive and 3 specific-differential queries. The descriptions of the query types are listed below:
@@
Opinion-based queries
Based on the core facts, we created 16 opinion-based queries consisting of 6 neutral, 5 confirmatory-original, and 5 confirmatory-alternate queries. We translated the queries into the target languages using translate-shell. The descriptions of the query types are listed below:
@@ 3.1.2 Preference Score
To calculate the preference scores for each document, we normalize cosine similarity scores using z-score for the documents retrieved by each query. In this paper, we use the mean preference score to measure documents grouped by a property, such as Foreign Language vs. Native Language document.
@@ 3.2 Generation
In the generation process, we provide two documents in different languages as the context for generating the answer. We need reference answers for each context separately to evaluate which context the model used as the source for generating the answer. We then compare the generated answer to each reference answer from both document sets (original and alternate) to determine which context is more represented in the generated answer.
@@ 3.2.1 Evaluation in English:
For the evaluation of generation, we use English reference answers only and translate the generated answers back to English using translate-shell444translate-shell can be viewed here. This approach is necessary due to the lack of high-quality, robust metrics for evaluating multilingual text in generation tasks Ahuja et al. (2023).
@@ 3.2.3 Information overlap (IO):
We use Information Overlap (IO) as our primary evaluation method. We use different strategies to evaluate the queries depending on the query type. For queries with subjective or lengthy answers—such as broad queries in the festival premise and all queries in the war premise—we use BERTScoreZhang* et al. (2020) with the microsoft/deberta-xlarge-mnli model, which is the best-performing model in WMT16 for English tasks. We use keyword matching for factual queries—all specific queries on the festival premises.
@@
To analyze which document-set (original Language_i or alternate language_j) the model preferred to generate the answer, we use source reference proportion as our measure. We compute the IO of the gold_answer in the original language with the generated answer, which we call original_overlap and compare that with the IO of the gold_answer in the alternate language with the generated answer (alternate_overlap). The desired outcome is for the source reference proportion to be Both, i.e. when both scores are above $\theta$ or Original_overlap = Alternate_overlap. The outcome is Original when Original_overlap is higher and Alternate when Alternate_score is higher. The outcome is None when both scores are below $\alpha$.
We do not consider $\alpha$ and $\theta$ when the IO method is a keyword match. For BertScore, the authors manually analyzed and set $\alpha$ to 0.55, where answers above this threshold contained partial similarity with the reference answer but did not cover all points, and $\theta$ to 0.7, where answers above this threshold contained most of the information from the reference answer.
@@ 4.1.1 RQ1r: Preference for Native Language
Overall, for the top-10 retrieved documents, models preferred the Native Language Document 68% of the time across both premises. As shown in Figure 4, there was a strong preference for Native Language document retrieval, with an average z-score of 1.03 for Native Language versus -0.25 for Foreign Language. The trend is consistent across all models. Further results are in Appx. LABEL:sec:_appendix-retrieval
@@ 4.1.2 RQ2r: Impact of query type on language preference
For factual queries, we did not observe significant differences across question types for both Native Language preference (Broad: M = 1.01, SE = 0.24; Specific-Contradiction: M = 0.97, SE = 0.21; Specific-Exclusive: M = 1.03, SE = 0.25; Specific-Differential: M = 0.90, SE = 0.22) and Foreign Language preference (Broad: M = -0.25, SE = 0.06; Specific-Contradiction: M = -0.24, SE = 0.05; Specific-Exclusive: M = -0.25, SE = 0.06; Specific-Differential: M = -0.22, SE = 0.05). The results are consistent across all models.
Similarly, for opinion-based queries, no significant differences were found across different question types for both Native Language preference (Neutral: M = 0.80, SE = 0.27; Confirmatory: M = 0.87, SE = 0.26) and Foreign Language Preference (Neutral: M = -0.20, SE = 0.06; Confirmatory: M = -0.21, SE = 0.06). Native Language documents are more likely to be retrieved by all models.

### v3

@@ Abstract
Although the multilingual capability of LLMs offers new opportunities to overcome the language barrier, do these capabilities translate into real-life scenarios where linguistic divide and knowledge conflicts between multilingual sources are known occurrences? In this paper, we studied LLM’s linguistic preference in a cross-language RAG-based information search setting. We found that LLMs displayed systemic bias towards information in the same language as the query language in both document retrieval and answer generation. Furthermore, in scenarios where no information is in the language of the query, LLMs prefer documents in high-resource languages during generation, potentially reinforcing the dominant views. Such bias exists for both factual and opinion-based queries. Our results highlight the linguistic divide within multilingual LLMs in information search systems. The seemingly beneficial multilingual capability of LLMs may backfire on information parity by reinforcing language-specific information cocoons or filter bubbles further marginalizing low-resource views.
@@
Figure 1: The figure shows users interacting with an LLM using 4 languages (en, ar, hi, zh). When the contexts are available in 3 (en,zh,hi) of the 4 languages, the queries in different languages resulted in drastically different answers for each language. We see 2 effects in this figure: 1) Information cocoon – information presented to the users of a particular language is from their own language and 2) Cultural Dominance – information asked from a lower resource language where the information is not available in, results in LLM choosing highest resource language as the information source leading to reinforcement of dominant views.
1 Background and Motivation
@@
With Retrieval Augmented Generation (RAG), Large Language Models (LLMs) can leverage external knowledge and respond to questions about ongoing events Lewis et al. (2021). More importantly, recent strides in LLM’s multilingual capabilities  Qin et al. (2025); Watts et al. (2024), supporting over 100 languages, offer new opportunities to break the language-specific information cocoon and provide more equitable information access for people with diverse linguistic preferences, cultural backgrounds, and geographical locations.
@@
However, achieving information parity is challenging because of the intrinsic complexity of how information is created and represented in different linguistic contexts. This complexity stems from a broad spectrum of cultural nuances, regional influences, and historical narratives that shape how information is consumed and communicated White (1990); Edelman (1985); Sittar et al. (2022); Lück et al. (2018). In addition, each language and cultural context also brings its own set of values and perspectives when interpreting an event of interest, which can significantly influence the representation of facts and narratives Foucault (1980).
@@
This challenge is exacerbated by the linguistic divide, the gap in representation across different languages li2024quantifying; Mehdad et al. (2012). Not all facts and viewpoints are presented equally. Dominant languages, such as English, have a disproportionate influence on content creation and dissemination Holborow (1996). Such divides lead to linguistic imperialism, where information in higher-resource languages is more accessible and frequently amplified, potentially overshadowing or distorting narratives in low-resource languagesPhillipson (2018), which can lead to mistrust in information seeking Arguedas et al. (2023).
@@
As a result, RAG systems need to retrieve from diverse multilingual sources and generate meaningful answers, even with conflicting knowledge across languages. For example, when an information seeker asks for an event of conflict or topic with multiple perspectives in different regions, like an ongoing war between two countries speaking distinct languages, LLMs should be able to present viewpoints and facts from both sides regardless of the query language.
@@
However, recent evidence indicates information disparity in LLMs’ behavior may stem from unequal language resources in current multilingual LLMsYu et al. (2022); Verma et al. (2022); Joshi et al. (2020) where pretraining data of models such as Llama and GPT-4 is still predominantly English Li et al. (2024b). As a result, individuals asking the same query in different languages may get different answers, Figure 1. This difference manifests in the quality of the responses Boughorbel and Hawasly (2023); Jin et al. (2024a), consistency Dong et al. (2024), and dispute resolution Li et al. (2024a).
Active research areas relevant to this domain include Cross-Language Information Retrieval (CLIR) and Multilingual IR (MLIR), which aim to retrieve documents from multilingual sourcesGrefenstette (1998); Carbonell et al. (1997); Guo et al. (2024); Zhang et al. (2020); Nair et al. (2022); Zhang et al. (2023); Yang et al. (2024); Parton et al. (2008); Yang et al. (1998). However, the aim of MLIR is to retrieve the most relevant documents regardless of language, without addressing conflicts or disparities of data across languages. These models struggle with information parity and knowledge conflicts, as we demonstrate in this paper. Likewise, Wu et al. (2024), highlighted similar challenges in low-resource CLIR settings.
Another research area focuses on knowledge conflicts and diverse perspectivesChen et al. (2019); Zhang et al. (2011); Chen et al. (2022); Cohen et al. (2024). Literature has studied three types of knowledge conflicts: inter-context, intra-memory, and context-memoryXu et al. (2024). However, there is little overlap between CLIR research and studies on knowledge conflicts. Moreover, they do not explore the linguistic divide’s effect on information disparity within multilingual LLMs. In RAG, biases may favor information retrieval and generation from certain languages especially when interacting with conflicting and contradictory information across different languages. Such biases threaten the goal of using multilingual LLMs for democratized global information access. If unaddressed, they may reinforce cultural dominance and create an information cocoon, alienating speakers of non-dominant languages.
This study uniquely aims to understand information parity in multilingual retrieval and generation individually where conflicting perspectives and facts are presented in different languages in an inter-context setting. Furthermore, we also aim to uncover the influence of diverse user query types motivated by the findings of Sharma et al. (2024), where users issued confirmatory queries while interacting with controversial information, which hasn’t been studied before. Specifically, we ask the following research questions,
@@ 2 Experiments
There are two types of LLM-powered search systems: Retrieval Augmented Generation (RAG) and Direct Generation. In this paper, we focus on RAG Lewis et al. (2021), where LLMs primarily rely on external knowledge presented in the context to answer user queries. We made this choice for the following reasons: 1) most popular LLMs do not provide access to their parametric memory, and 2) most popular information systems, such as Bing Copilot, Perplexity.ai, etc., use RAG to keep their search system updated with the most recent information.
@@
RAG consists of two phases: retrieval and generation. In the retrieval phase, the system retrieves a ranked list of documents based on cosine similarity scores between query embedding and document embedding. In the generation phase, the top documents are used as LLM’s context to generate the response. As studied previously when dealing with conflicts generation is impacted by performance of retrievalChen et al. (2022) and hence in this study, we study the two phases independently.
Need for Synthetic Data
LLM’s parametric memory and biases distinctly influence the resultsJin et al. (2024b); Xie et al. (2024), making it essential to distinguish between parametric and in-context non-parametric memory effects. In addition, real-world controversies may activate guardrails in high-resource languages such as English, which can confound our results Deng et al. (2024). To mitigate these influences, we developed a synthetic dataset featuring fictional entities absent from the LLM’s pre-training data and guardrails.
@@ 2.1 Multilingual Fictional Dataset
Figure 2: The figure outlines our fictional dataset creation: core facts were crafted manually, then expanded using GPT-4-turbo OpenAI (2024) without contradictions. Different subsets of these expanded facts were used to generate News and Story documents using the LLM. This process yielded both original and alternate document sets. We then translated these documents into target languages using Google Translate via translate-shell. Prompt details are in the Appendix.
The overarching goal of our dataset is to emulate a real-world multilingual information environment where different, sometimes conflicting, facts and perspectives are represented in different languages. The dataset consists of 170 documents in 5 languages that were later used in our experiment to answer user queries. The dataset creation steps and manipulation checks are detailed in Appx. A.2.
As highlighted in Fig. 3 and Appx. A.1, we support knowledge conflicts manifesting either as: 1) discrepancies in facts where same event in different languages contains different or conflicting facts due to differences in cultural and historical narratives or 2) conflicting perspectives and opinions where complex geopolitical issues, may have documents describe the same event with different perspectives and opinions in different languages.
@@
We selected English (en), Hindi (hi), German (de), Arabic (ar), and Chinese (zh) as our target languages. These languages represent 4 different scripts, different cultures, and a mix of higher-resource (en, de, zh) and lower-resource languages (ar, hi). These languages account for around 4 billion active speakers around the world.
@@ 2.1.1 Premise setup
The festival premise is designed for factual queries, which contain information about a fictional festival. There are two versions: original and alternate – each with some shared, contradictory, and exclusive facts. Here is an example with contradictory facts,
@@
The war premise is designed for opinion-based queries. It contains information about a fictional war taking place between two fictional regions speaking different languages. This premise is designed to represent an information environment with diverse opinions and perspectives. Similarly, there are two document sets in the war: original and alternate – each taking the perspective of one side in the war. Example perspectives are,
@@ 2.2 Experiment Design
We queried popular multilingual retrieval models and LLMs with different query types to analyze the preference of models when presented with multilingual information conflicts.
@@ 2.2.1 Diverse Information Queries
We curated a diverse set of queries to mimic real-world multilingual information-seeking behaviors. We covered two types of information queries: factual and opinion-based. Each query is translated from English to Hindi(hi), Chinese(zh), Arabic(ar), and German(de) using translate-shell.
@@
Based on the core facts, we created 27 factual queries consisting of 6 broad, 9 specific-contradiction, 9 specific-exclusive and 3 specific-differential queries:
@@
Opinion queries
We created 16 opinion-based queries consisting of 6 neutral, 5 confirmatory-original, and 5 confirmatory-alternate queries:
@@ 3.1.2 Preference Score
To calculate preference scores for each document, we normalize cosine similarity scores using z-score for the documents retrieved by each query. In this paper, we use the mean preference score to measure documents grouped by a property, such as Foreign Language vs. Native Language documents.
@@ 3.2 Generation
Two documents in different languages serve as the context for generating the answer. Reference answers for each context are needed to evaluate which document the model used as the source. We then compare the generated answer to each reference answer from both document sets (original and alternate) to identify which context is more represented in the generated answer.
@@ 3.2.1 Evaluation in English:
For the evaluation of generation, we use English reference answers only and translate the generated answers back to English using translate-shell111translate-shell can be viewed here. This approach is necessary due to the lack of high-quality, robust metrics for evaluating multilingual text in generation tasks Ahuja et al. (2023).
@@ 3.2.3 Information overlap (IO):
We use Information Overlap (IO) as our primary evaluation method. We use different strategies to evaluate the queries depending on the query type. For queries with subjective or lengthy answers—such as broad queries in the festival premise and all queries in the war premise — we use BERTScoreZhang* et al. (2020) with the microsoft/deberta-xlarge-mnli model (best-performing model in WMT16 for English tasks). We use keyword matching for factual queries—all specific queries on the festival premises.
@@
We use source reference proportion to determine which document (original vs alternate) the model preferred to generate the answer. We compare the IO between the alternate_overlap generated answer and the gold answer from the original (original_overlap) and alternate document(alternate_overlap). The desired outcome is "Both", i.e. when both scores are above $\theta$ or Original_overlap = Alternate_overlap. The outcome is Original when Original_overlap is higher and Alternate when Alternate_score is higher. The outcome is None when both scores are below $\alpha$. For keyword match, $\alpha$ and $\theta$ are not applied. For BERTScore, after manual analysis, $\alpha$ is set to 0.55 for partial similarity, and $\theta$ is set to 0.7 for when most information from the reference is covered.
@@ 4.1.1 RQ1r: Preference for Native Language
Looking at top-10 retrieved documents, models preferred the Native Language Document 68% of the time across both premises. As shown in Figure 4, there was a strong preference for Native Language document retrieval, with an average z-score of 1.03 for Native Language versus -0.25 for Foreign Language. The trend is consistent across all models. Further results are in Appx. LABEL:sec:_appendix-retrieval
@@ 4.1.2 RQ2r: Impact of query type on language preference
For factual queries, we did not observe differences across question types for both Native Language preference (Broad: M = 1.01, SE = 0.24; Specific-Contradiction: M = 0.97, SE = 0.21; Specific-Exclusive: M = 1.03, SE = 0.25; Specific-Differential: M = 0.90, SE = 0.22) and Foreign Language preference (Broad: M = -0.25, SE = 0.06; Specific-Contradiction: M = -0.24, SE = 0.05; Specific-Exclusive: M = -0.25, SE = 0.06; Specific-Differential: M = -0.22, SE = 0.05). The results are consistent across all models.
Similarly, for opinion-based queries, no differences were found across different question types for both Native Language preference (Neutral: M = 0.80, SE = 0.27; Confirmatory: M = 0.87, SE = 0.26) and Foreign Language Preference (Neutral: M = -0.20, SE = 0.06; Confirmatory: M = -0.21, SE = 0.06). Native Language documents are more likely to be retrieved by all models.